OpenStudy (chrisplusian):

Show that there are no vectors "u" and "v" in R^3 such that ||u||=1, ||v||=2, and =3.

OpenStudy (loser66):

Suppose yes, then \(cos (u, v) =\dfrac{<u, v>}{||u||||v||}= \dfrac{3}{2}\) does it make sense? You give out the conclusion.

OpenStudy (chrisplusian):

I honestly don't follow that way.

OpenStudy (loser66):

why not? \(-1\leq cos (\theta) \leq 1\) That is the restriction of a cos function, and the result is out of the range of cos. Hence the "supposing" does not hold .

OpenStudy (loser66):

I use contradiction method and it is the easiest way to use all of the given information. But if you don't get it. I am sorry.

OpenStudy (chrisplusian):

Thank you for the attempt, and I don't doubt what you have done. I just think there has to be a different way to show it. Can you think of a different way to do it?

OpenStudy (loser66):

@oldrin.bataku

OpenStudy (anonymous):

$$\langle v,u\rangle=\|u\|\|v\|\cos\theta=2\cos\theta$$however \(\cos\theta\) has a maximum value of \(1\) so \(2\cos\theta\le 2\), yet \(3\not\le 2\) so there can be no such vectors

OpenStudy (anonymous):

geometrically, we have that \(\langle v,u\rangle\le\|v\|\|w\|\), and \(3\not\le2\)

OpenStudy (chrisplusian):

Is there a way to expand the dot product to show this?